You have 15 coins in your pocket that are either quarters or nickels. They total $2.75. How many of each coin do you have?
Answer by checkley75(3666)

n+q=15
n=15-q
5n+25q=2.75
5(15-q)+25q=2.75
75-5q+25q=2.75
20q=2.75-75
20q=200
q=200/20
q=10 answer for number of quarters thus
n+10-15
n=15-10
n=5 number of nickels
proof
5*.05+10*.25=2.75
.25+2.50=2.75
2.75=2.75